Can you please help me with these questions?

Use the wave equation to find the speed of a wave given by 0.5 y(x, t) = (2.00 mm) | (20 m-1 )x-(4.0 s- Use the wave equation to find the speed of a wave given in terms of the general function hx, t): y(x, t) = (4,00 mm) hl (30 m-1)x + (6.0 s-1
Answer #1

from general equation

y(x,t)= Asin(kx-wt)

w=kv

v=k/w

1. v=20/4=5m/s

2. v=30/6=5m/s